# Autonomous Issue Resolver: Towards Zero-Touch Code Maintenance

### v1

@@ Abstract
Recent advances in Large Language Models have revolutionized function-level code generation; however, repository-scale Automated Program Repair (APR) remains a significant challenge. Current approaches typically employ a control-centric paradigm, forcing agents to navigate complex directory structures and irrelevant control logic. In this paper, we propose a paradigm shift from the standard Code Property Graphs (CPGs) to the concept of Data Transformation Graph (DTG) that inverts the topology by modeling data states as nodes and functions as edges, enabling agents to trace logic defects through data lineage rather than control flow. We introduce a multi-agent framework that reconciles data integrity navigation with control flow logic. Our theoretical analysis and case studies demonstrate that this approach resolves the ”Semantic Trap” inherent in standard RAG systems in modern coding agents. We provide a comprehensive implementation in the form of Autonomous Issue Resolver (AIR), a self-improvement system for zero-touch code maintenance that utilizes neuro-symbolic reasoning and uses the DTG structure for scalable logic repair. Our approach directly addresses the core limitations of current AI code-assistant tools and tackles the critical need for a more robust foundation for our increasingly software-dependent world.
@@ 1 Introduction
The early days of software developed by smaller teams are long gone. Today’s software products are huge, often comprising millions of lines of code and hundreds of dependencies. No single human, nor even large teams, has a comprehensive understanding of such systems. This is the same as trying to build a skyscraper with bare hands. Current development methodologies attempt to manage this complexity by breaking down the problems, but this merely distributes the potential failures among developers. This is because individual developers, working on isolated components, still introduce bugs, inconsistencies, and architectural drift. The consequences are significant: costly delays, security vulnerabilities, system outages, and a drag on innovation. Industry reports consistently highlight that debugging and maintenance consume a staggering 50-70% of development resources [1].
The path to truly reliable software lies in systematically reducing the dependence on manual intervention for repetitive, error-prone, and complex analytical tasks. AI-based automation offers the potential for consistency, speed, and depth of analysis that exceed human capabilities. By automating many stages of the software development lifecycle (SDLC), from code generation and testing to debugging and refactoring, we can mitigate human error, enforce best practices, and achieve unprecedented levels of quality assurance.
@@ 3 Related Work
The introduction and further development of Transformer-based models trained on large code corpora has revolutionized software engineering. Early breakthroughs like Codex [4] and AlphaCode [5] demonstrated that models could synthesize functional code from natural language descriptions. These foundational models power a generation of interactive assistants, including GitHub Copilot [6], Amazon CodeWhisperer [7], and Cursor [8]. While effective for local, ”next-block” completion [2], these tools primarily operate as ”copilots,” relying on the human developer to manage the broader repository context and verify correctness.
@@
Agentic frameworks such as ReAct [9] and Reflexion [10], allowed LLMs to plan, execute tools, and observe outputs. State-of-the-art systems like SWE-agent [11] and OpenHands [12] utilize these patterns to interact with repositories via shell commands and file editors, and currently often used as solid baselines. Similarly, AutoCodeRover [13] employs abstract syntax tree (AST) analysis to improve fault localization.
@@
However, these agents largely treat the repository as a file system to be searched and read, a ”file-first” perspective that suffers from inefficiency when bugs span multiple abstraction layers. Recent evaluations on benchmarks like SWE-bench [14, 15] indicate that while agents can solve isolated tasks, they struggle with long-horizon reasoning required for enterprise issues [16]. Claude Code [17] and LocAgent [18] attempt to mitigate this by improving retrieval and localization, yet they still fundamentally operate on text-based or partially structured representations that do not explicitly model data lineage.
@@
Since the search over codebase is essentially a search over unstructured data, it was natural to employ the graph-based methods for localization of relevant snippets of code. The Code Property Graph (CPG) [19] unified ASTs, control-flow graphs (CFGs), and program dependence graphs (PDGs) into a single structure, primarily for vulnerability discovery.
@@
Recent efforts have attempted to integrate such graphs with LLMs to improve repository understanding. CodexGraph [20] and RepoGraph [21] utilize graph databases to enhance Retrieval-Augmented Generation (RAG) by capturing dependencies between files and functions. The Code Graph Model (CGM) [22] similarly integrates graph structures to aid in logical reasoning. However, these approaches often retain the ”representation bloat” of standard CPGs or focus on improving retrieval recall rather than navigation that mimics the ”operational reasoning”. Unlike these control-centric models, our approach utilizes a Data-First Transformation Graph (DTG), which compresses the topology to focus on data states and transformations, aligning somehow our approach with the ”flow-based” mental models used in high-performance computing optimization.
@@
Although most current agents rely on prompt engineering and in-context learning, Reinforcement Learning (RL) has shown promise in discovering novel strategies in complex domains. DeepMind’s work on AlphaZero [23] and subsequently AlphaDev [24] demonstrated that RL agents could discover algorithms that outperform human-written implementations. In the domain of program repair, early learning-based systems like Getafix [25] mined fix patterns from static analysis. More recently, AlphaEvolve [26] pushes this further by evolving code structures. Our work incorporates an RL-driven control policy to navigate the DTG, distinguishing AIR from purely prompt-driven agents by allowing it to learn stopping conditions and navigation strategies that minimize regression risks. We are using a simple DQN-based online learning. Note that RL is not a key element of our algorithm, and it can be replaced by simple Monte Carlo Tree Search style graph traversing.
@@ 4.1 A data-first graph representation
Figure 4.1: The partial visualization of the multi-dimensional Data Transformation Graph for a snippet of code. Nodes contains full description of processed data. Edges are effectively transformation functions (sanitize, handle)
@@ 5.1 Code property graphs and representation bloat
The Code Property Graph (CPG), popularized by tools such as Joern [28], unifies three classic program representations in a single multi-graph: abstract syntax trees (ASTs), control-flow graphs (CFGs), and program dependence graphs (PDGs). In principle, CPGs contain all information needed for both security analysis and program repair.
@@ 5.4.1 Formal definition
  • name: A unique identifier, e.g., module.func.var.
  • kind: Class of data, e.g. is it the argument of function, a global variable, or constant, etc.
@@ 6.1 Agent environment
7 High-Level Architecture (Product)
The design of the competitive AI tool for autonomous code development requires careful consideration of its integration into existing developer workflows and ecosystems. This section details the rationale behind our core architectural decisions: adopting a frontless application model and pursuing deep integration with Git service platforms.
7.1 Why We Designed a Frontless Application
The primary goal of our AI product is to enhance and automate the software development lifecycle, particularly within Continuous Integration and Continuous Deployment (CI/CD) pipelines, to accelerate development efforts.
Modern software development heavily relies on automated CI/CD pipelines. These pipelines are inherently process-driven, triggered by events like code commits or pull requests, and often operate without direct human interaction. Unlike tools designed for interactive use (such as IDE plugins or code editors), our product aims to perform complex tasks with minimal or zero real-time human guidance once configured. The ”interface” becomes the configuration files and the standard platform mechanisms for feedback (PR comments, commit statuses, logs). This aligns with the concept of ”Invisible Computing” or ”Calm Technology,” [30] where technology recedes into the background, providing value without demanding constant attention.
Headless systems are often easier to scale and integrate via APIs and webhooks. Our AI can be deployed as a service that listens for events from multiple repositories or projects simultaneously, processing tasks asynchronously. This architecture is well-suited for organizational-level adoption where the tool needs to operate across numerous CI/CD pipelines.
7.2 On the Importance of Platform
While a frontless architecture defines how the user interacts (or does not directly interact) with our tool, where the tool operates is equally critical. Deep integration within established Git service platforms such as GitHub and GitLab is our primary technical and business strategy.
Platform integration allows our AI to become a native participant in the core developer workflow – the Pull/Merge Request process. It can be triggered automatically by platform events (PR opened, commit pushed, comment posted), perform its analysis and code generation, and report results directly back into the platform interface (e.g., posting PR comments with suggestions, pushing code changes to the PR branch, updating status checks). Our tool operates autonomously and asynchronously, typically after the initial code has been written (Figure 7.3)
There is another, more subtle reason to build a platform app rather than a standalone one: platforms like GitHub and GitLab offer rich ecosystems. Integration allows our tool to potentially interact with other tools and services operating within the same platform. This creates opportunities for more powerful, composite automation, and sometimes this alone is valued much more than just a yet another fancy tool for code generation.
Figure 7.3: Our product leverages the existing GitHub platform and workflows (Issues, Branches, PRs). There is no need for developers to learn new tools or interfaces. The Bot acts like another member of the team, communicating and contributing via established mechanisms. Such approach does not break existing automation. Note, that human developers can still review the Bot’s code changes directly within the PR interface, leave comments, or request modifications if needed. This step maintains human oversight where desired.
7.3 Distinction from Interactive Code Assistants (e.g., GitHub Copilot)
Interactive assistants like GitHub Copilot primarily focus on augmenting the developer during the act of writing code within the IDE. They operate synchronously, responding to immediate prompts and context within the editor. Our tool operates autonomously and asynchronously, typically after the initial code has been written, focusing on broader, process-oriented tasks within the development life-cycle (e.g., automated refactoring based on PR feedback, generating tests for a new feature branch, ensuring code adheres to project-wide patterns).
Copilot assists the individual developer’s coding speed; our tool automates team and process workflows, operating at the level of commits, branches, and pull requests within the platform.
@@ Appendix B AI-assisted software tools
Table B.2 presents a comparison of market tier, interaction modality, and primary capabilities of today’s AI coding tools. While ”Professional” tier tools (e.g. Copilot, Cursor) dominate the interactive space via IDE integration, they operate primarily on a ”Human-in-the-Loop” synchronous model and hardly can be considered a serious AI technology. ”Specialized” agents (AlphaCode) excel at algorithmic synthesis but lack repository awareness. AIR is distinct in its focus on asynchronous, autonomous maintenance, targeting the ”Zero-Touch” quadrant currently unoccupied by major commercial providers.

### v2

@@ Abstract
Recent advances in Large Language Models have revolutionized function-level code generation; however, repository-scale Automated Program Repair (APR) remains a significant challenge. Current approaches typically employ a control-centric paradigm, forcing agents to navigate complex directory structures and irrelevant control logic. In this paper, we propose a paradigm shift from the standard Code Property Graphs (CPGs) to the concept of Data Transformation Graph (DTG) that inverts the topology by modeling data states as nodes and functions as edges, enabling agents to trace logic defects through data lineage rather than control flow. We introduce a multi-agent framework leveraging a Dual-Graph Consensus mechanism that reconciles data integrity navigation with control flow logic. Our theoretical analysis and case studies demonstrate that this approach resolves the ”Semantic Trap” inherent in standard RAG systems in modern coding agents. We provide a comprehensive implementation in the form of Autonomous Issue Resolver (AIR), a self-improvement system for zero-touch code maintenance that utilizes neuro-symbolic reasoning and uses the DTG structure for scalable logic repair. Our approach directly addresses the core limitations of current AI code-assistant tools and tackles the critical need for a more robust foundation for our increasingly software-dependent world.
@@ 1 Introduction
The early days of software developed by smaller teams are long gone. Today’s software products are huge, often comprising millions of lines of code and hundreds of dependencies. No single human, nor even large teams, has a comprehensive understanding of such systems. This is the same as trying to build a skyscraper with bare hands. Current development methodologies attempt to manage this complexity by breaking down the problems, but this merely distributes the potential failures among developers. This is because individual developers, working on isolated components, still introduce bugs, inconsistencies, and architectural drift. The consequences are significant: costly delays, security vulnerabilities, system outages, and a drag on innovation. Industry reports consistently highlight that debugging and maintenance consume a staggering 50-70% of development resources [1]. This ”maintenance tax” prevents organizations from focusing on new feature development, trapping them in a cycle of reactive patching.
The path to truly reliable software lies in systematically reducing the dependence on manual intervention for repetitive, error-prone, and complex analytical tasks. AI-based automation offers the potential for consistency, speed, and depth of analysis that exceed human capabilities. By automating many stages of the software development life-cycle (SDLC), from code generation and testing to debugging and refactoring, we can mitigate human error, enforce best practices, and achieve unprecedented levels of quality assurance.
@@ 3 Related Work
The field of LLM-based software engineering has exploded between 2024 and 2026, characterized by a dichotomy between ”Agentic” workflows (complex, multi-step reasoning) and ”Agentless” approaches (simplified, search-and-repair pipelines). The introduction and further development of Transformer-based models trained on large code corpora has revolutionized software engineering. Early breakthroughs like Codex [4] and AlphaCode [5] demonstrated that models could synthesize functional code from natural language descriptions. These foundational models power a generation of interactive assistants, including GitHub Copilot [6], Amazon CodeWhisperer [7], and Cursor [8]. Although effective for local ”next-block” completion [2], these tools primarily operate as ”copilots,” relying on the human developer to manage the broader repository context and verify correctness.
@@
Agentic frameworks such as ReAct [9] and Reflexion [10], allowed LLMs to plan, execute tools, and observe results. State-of-the-art systems like SWE-agent [11] and OpenHands [12] utilize these patterns to interact with repositories via shell commands and file editors, and are currently often used as solid baselines. Similarly, AutoCodeRover [13] employs abstract syntax tree (AST) analysis to improve fault localization.
@@
However, these agents largely treat the repository as a file system to be searched and read, a ”file-first” perspective that suffers from inefficiency when bugs span multiple abstraction layers. Recent evaluations on benchmarks such as SWE-bench [14, 15] indicate that while agents can solve isolated tasks, they struggle with long-horizon reasoning required for enterprise issues [16]. Claude Code [17] and LocAgent [18] attempt to mitigate this by improving retrieval and localization, yet they still fundamentally operate on text-based or partially structured representations that do not explicitly model data lineage.
@@
Since the search over the codebase is essentially a search over unstructured data, it was natural to employ the graph-based methods for localization of relevant snippets of code. The Code Property Graph (CPG) [19] unified ASTs, control-flow graphs (CFGs), and program dependence graphs (PDGs) into a single structure, primarily for vulnerability discovery.
@@
Recent efforts have attempted to integrate such graphs with LLMs to improve repository understanding. CodexGraph [20] and RepoGraph [21], [22] utilize graph databases to enhance Retrieval-Augmented Generation (RAG) by capturing dependencies between files and functions. The Code Graph Model (CGM) [23] similarly integrates graph structures to aid in logical reasoning. However, these approaches often retain the ”representation bloat” of standard CPGs or focus on improving retrieval recall rather than navigation that mimics the ”operational reasoning”. Unlike these control-centric models, our approach utilizes a Data-First Transformation Graph (DTG), which compresses the topology to focus on data states and transformations, somehow aligning our approach with the ”flow-based” mental models used in high-performance computing optimization.
@@
Although most current agents rely on prompt engineering and in-context learning, Reinforcement Learning (RL) has shown promise in discovering novel strategies in complex domains. DeepMind’s work on AlphaZero [24] and subsequently on AlphaDev [25] demonstrated that RL agents could discover algorithms that outperform human-written implementations. In the domain of program repair, early learning-based systems such as Getafix [26] mined fix patterns from static analysis. More recently, AlphaEvolve [27] pushed this further by evolving code structures. Our work incorporates an RL-driven control policy to navigate the DTG, distinguishing AIR from purely prompt-driven agents by allowing it to learn stopping conditions and navigation strategies that minimize regression risks. We are using a simple DQN-based online learning. Note that RL is not a key element of our algorithm, and it can be replaced by simple Monte Carlo Tree Search style graph traversing.
@@ 4.1 A data-first graph representation
Figure 4.1: The partial visualization of the multi-dimensional Data Transformation Graph for a snippet of code. Nodes contains full description of processed data. Edges are effectively transformation functions (sanitize, handle). The arrows demonstrate the ”causal path” of the variable id as it traverses from the process_transactions scope (blue) into the sanitize scope (yellow) and returns as a new entity. This structure achieves Semantic Compression, abstracting away syntactic noise (AST nodes) and enabling the agent to navigate logic defects purely through data lineage.
@@ 5.1 Code property graphs and representation bloat
The Code Property Graph (CPG), popularized by tools such as Joern [29] and PyXray [30], unifies three classic program representations in a single multi-graph: abstract syntax trees (ASTs), control-flow graphs (CFGs), and program dependence graphs (PDGs). In principle, CPGs contain all information needed for both security analysis and program repair.
@@ 5.4.1 Formal definition
  • full_name: A unique identifier, e.g., module.func.var.
  • kind: Class of data, e.g. is it the argument of a function, a global variable, or constant, etc.
@@ 6 Graph construction pipeline
A key theoretical advantage of the DTG is Semantic Compression. A standard CPG for a function def foo(x): if x>0: return x+1 might contain 20+ nodes (AST nodes for If, Compare, Return, BinOp, etc.). [19] The DTG compresses this entire subgraph into a single edge: $v_{in}\xrightarrow{foo[x>0]}v_{out}$ This reduces the graph size by orders of magnitude (approx. 100x), enabling the agent to load the ”skeleton” of the entire repository into its working memory, solving the context window bottleneck.
Pure data-flow graphs lose control-flow information (branches). We reintroduce this via Guarded Edges. An edge is annotated with a guard $\gamma$ derived from the CFG.
$v_{j}=T(v_{i})\iff\gamma(v_{control})=\text{True}$
This allows the agent to reason: ”This transformation only happens if the is_verified flag is set,” maintaining the necessary logic for bug reproduction without the bloat of the full CFG.
@@ Appendix B AI-assisted software tools
Table B.2 presents a comparison of market tier, interaction modality, and primary capabilities of today’s AI coding tools. While ”Professional” tier tools (e.g. Copilot, Cursor) dominate the interactive space via IDE integration, they operate primarily on a ”Human-in-the-Loop” synchronous model and hardly can be considered a serious AI technology. ”Specialized” agents (AlphaCode) excel at algorithmic synthesis but lack repository awareness. AIR is distinct in its focus on asynchronous, autonomous maintenance, targeting the ”Zero-Touch” quadrant.